### v1

@@ Abstract
Data Fusion, Foundational Models, Embeddings, Multimodal Data
@@ 1 Introduction
In this era, vast amounts of data and information are generated in different settings, fields, formats, and modalities. According to the International Data Corporation (IDC), 7.5ZB are expected to be generated every year [1], and the ability to process this information and extract knowledge has never been more important [2, 3]. This data can come from diverse sources such as wearable devices, laboratory tests, Electronic Health Records (EHR), among many others, and modalities, such as medical images, clinical notes, and vital signs. The integration of data from different modalities is on the rise in Artificial Intelligence (AI) fields, such as Machine Learning (ML) and Deep Learning (DL), known as multimodal data fusion. This research area, with applications integrating areas from natural language processing to computer vision and beyond, has driven applications in healthcare [4, 5], integrating EHRs with medical images [6, 7, 8, 9] or signals from wearable devices [10, 11]. Multimodal fusion has also led to applications such as autonomous driving [12, 13], environmental sciences applications combining different sensors and satellite data[14, 15], as well as many other Internet of Things (IoT) applications and system improvements[16, 17, 18].
@@
Recent advances in the DL field, mainly driven by the proposal of the transformer architecture [19], have seen breakthroughs in the analysis of different modalities, such as text with models like BERT [20] or GPT [21] or images with models like Vision Transformers [22]. These models have been a revolution also in specific fields such as medicine with impressive performances in medical applications such as medical tests or diagnosis [23, 24, 25]. However, some of the limitations of working with these data modalities and these specialized DL models are, the high dimensionality of the data, the heterogeneity of the diverse data modalities and formats, and the amount of data and computational resources required to train robust models for these specific modalities [26, 27]. Embeddings are low-dimensional numerical vectors learned during DL model training that preserve the original data’s most relevant features. Embeddings provide a common format that can solve data extraction and model development in environments where computational resources are scarce. Furthermore, their simplicity gives embeddings a possible way to unify the different data modalities and an efficient and generalizable solution for building multimodal models.
@@
Since embedding is proposed as a potential solution to data heterogeneity, and computational resource constraints, the DL models used to extract these embeddings must be extremely robust models trained on large amounts of data. In this context, we discuss open-source foundational models, such as Dino v2 [28] for general images or Llama 2 [29] for general text. Foundational models may also be available for specific tasks such as clinical notes with examples like Med-Bert [30], Bio-Bert [31], or medical images like Retfound for retinal images [32].
@@
However, having data and making a prediction is not enough and can cause harm if not done correctly. Adopting robust and reproducible practice is the most important characteristic of any data-related process [33]. Using AI techniques irresponsibly will lead to poor model performance in bias [34, 35, 36], robustness, and fairness [37, 38]. Acknowledging and addressing potential biases inherent in the data sources and models employed is pivotal.
@@
There are already previous frameworks for data fusion, starting with the best-known Joint Directors of Laboratories (JDL) [39] and its multiple updates that led to the Data Fusion Information Group (DFIG) [40] model with its most recent version that integrates AI techniques [41]. However, these models do not leverage foundational models and DL nor incorporate best practices for reducing bias and improving the model’s robustness, and have a focus only on applications such as industry and the military, leaving out use cases in environments such as healthcare.
@@
In this paper, we delve into a foundational approach for multimodal data fusion, centering the framework on the use of foundational models, vector embeddings, and good practices for data mining. Our work draws inspiration from established frameworks such as the JDL [39] and DFIG [40]. We emphasize the role of embeddings in providing flexibility and improving efficiency, robustness, and fairness, which is even more crucial in low-resource settings where errors and bias are relatively more impactful [42]. We also showed the model’s performance in three healthcare use cases: diabetic retinopathy diagnosis using fundus retinal images and EHR data, domestic violence prediction using open data from satellite images and internet data such as online news, and medical diagnosis and gender extraction using radiological images and clinical notes.
@@
  • Introduction of a Novel Data Fusion Model: We have developed an innovative process model for multimodal data fusion, rooted in the principles of foundational models and embeddings. This model is designed to efficiently address the challenges of big data’s high dimensionality and heterogeneity, making it particularly suitable for complex environments like healthcare.
@@ 2 Background
Some improvements were proposed during the following years, such as introducing adaptive generalized closed-loop data fusion processes and the perceptual reasoning system [45]. Eventually, a second revision of the JDL model (JDL II) to align with Department of Defense (DoD) priorities, improve quality control, reliability, and consistency, and empathize with the need for co-processing of inferencing processes was proposed in 2004 [46]. Also in that year, due to the shortcomings of the model, mainly related to the lack of clarity of human participation and feedback, the DFIG model was created, adding two levels (Level 5 - User Refinement, Level 6 - Mission Management) to the JDL model [40]. However, other adaptations and updates were proposed to the JDL model during the following years, mainly focusing on the optimization and reliability of the model by proposing communication across the JDL levels [47] or implementing different techniques such as game theory [48, 49], fuzzy logic [50], ontologies [51] and other mathematical modeling approaches [52, 53]. Other revisions focused on the fusion, impact, and communication between hard data (machine or environmental data) and soft data (human data). Changes in specific levels such as Levels 1 and 2 incorporating the human in the fusion and preprocessing levels [54], subsequently in Levels 4 and 5 through hybrid-sensing and hybrid-cognition [55], or even the incorporating a management network in parallel to the JDL model [56, 57] were proposed.
@@
The proposed Data Fusion for Data Mining (DF-DM) model is a foundational process model for multimodal data fusion based on DFGI and CRISP-DM. The following sections will introduce the most up-to-date data fusion model (DFGI) with the adaptation for ML and AI [41], integrate the Cross-Industry Standard Process for Data Mining (CRISP-DM) process model [59], and present some modifications to implement foundational models and embeddings in the DFGI model. Finally we’ll provide three use cases to show the flexibility and adaptability of our model, as well as usability.
@@ 2.2 The Integration of the Cross-Industry Standard Process for Data Mining (CRISP-DM)
  • Many modalities, many costs: The use of multimodal data, although one of the most efficient ways to integrate and take advantage of data, also becomes very expensive when discussing high-dimensional data such as text or images. Given this, we propose using newer techniques and models for dimensionality reduction, such as extracting embeddings from foundational models or using foundational models directly for zero-shot learning tasks.
@@
In the era of AI-driven analytics, the integration of data mining and data fusion stands as a paramount challenge, primarily due to the heterogeneity of data and the constraints of computational resources. To address these challenges, we introduce an innovative methodology within our Data Fusion for Data Mining (DF-DM) model (Figure LABEL:2), which extends the existing DFGI model by integrating the model with the best practices of the CRISP-DM process model. Additionally our model leverages the use of foundational models and vector embeddings simplifies the processing of diverse data types but also significantly reduces the computational overhead, making advanced data analytics more accessible and efficient.
@@
The rationale behind the DF-DM model’s design is twofold. Firstly, by incorporating the CRISP-DM process, we aim to adhere to a proven, structured approach for data mining that emphasizes business understanding, data understanding, bias, and a cyclical process for model refinement. This ensures that our model is not only technically sound but also closely aligned with practical, real-world applications. Secondly, the decision to utilize embeddings and foundational models stems from their ability to reduce high-dimensional data into more manageable, lower-dimensional vectors. This significantly alleviates computational demands and facilitates more effective integration of diverse data types, crucial for addressing the complex challenges of multimodal data fusion in healthcare and beyond. Before delving into the methodology, it is pertinent to clarify two pivotal concepts:
  • Foundational models: Foundational models are at the heart of our methodology. These are extensive, pre-trained models that have been developed to capture a wide array of features across specific data modalities, such as text or images. By understanding the general characteristics inherent to each modality, these models can generate useful representations of new, unseen data. This ability is crucial for extracting relevant features without necessitating additional, computationally expensive training phases for each new dataset.
  • Embeddings: Embeddings are also a pivotal role in our methodology, transforming high-dimensional data into a more manageable, lower-dimensional space. Formally, given an input data point $x\in\mathbb{R}^{n}$, where $n$ is the dimensionality of the raw data in a high-dimensional space. We seek to find its embedding $E\in\mathbb{R}^{d}$ in a d-dimensional vector space where $d$ is the dimension of the embedding vector and $d<<n$. The embedding is generated given the transformation $E=f(x)$, where $f$ denotes the transformation function derived from the foundational model. This process significantly reduces dimensionality, with $d<<dim(X)$, facilitating more efficient data analysis.
@@ 3.1 Level 0 - Data Assessment:
  • Data preprocessing: Data preprocessing involves cleaning, transformation, normalization, encoding, reduction, and formatting, making it a critical step in the overall DF-DM model. In this step, as shown in Figure 3, using embeddings and foundational models is crucial for generating latent representations and extracting meaningful information from high-dimensional data such as images or text. Techniques such as zero-shot learning or embedding extraction are the key elements in this approach to extract features in low-resource settings. This approach significantly reduces computational costs and storage requirements and creates a unified format that facilitates multimodal data fusion.
@@ 3.2 Level 1 - Object Assessment
In Level 1, since the data received from Level 0 are embeddings or features extracted from the foundational models, the data is already in a lower dimensionality and a common format, facilitating different ML and AI techniques. As seen in Figure 3, Option 2, using embeddings, reduces the complexity and dimensionality of the data, making it easy and inexpensive to do the ML modeling because we can use simple models such as linear models. The data resulting from embedding extraction is methodically divided based on data modality (e.g., text, images). This separation enables targeted model training for each data type, enhancing specificity and accuracy. Integration of these segmented models is achieved through a comprehensive analysis framework, allowing for nuanced insights across modalities. For example, using the data from Level 0 in a CSV format due to the embedding extraction, the CSV file can be split for each modality to train a predictive model per modality. ML models can be trained to separately measure the impact of clinical notes, medical images, and tabular data in a given condition.
@@ 3.8 Level 7 - Bias Assessment
Figure 3: Illustrative Framework for Utilizing Foundational Models in Various Tasks. The figure assumes an initial foundational model for a general task and 3 different options. Option 1 is Zero-shot learning using the foundational model directly for a downstream task. Option 2 suggests the use of embedding, where embeddings of the original data are extracted and used for downstream task training. Option 3 means fine-tuning the full model for a specific task. The resulting model can also be used for embedding extraction.
@@ 4.2 Model Bias
Model bias can occur when the model’s architecture or training data introduces inherent bias. For example, if the foundational models used for embedding extraction have been trained on biased datasets, this bias can propagate into the DF-DM model. To address model bias, we suggest considering the following:
  • Bias Analysis of Foundational Models: Before embedding extraction, thoroughly analyze the foundational models to identify and mitigate potential bias. Analyze the current model available and the best option for your use case. Characteristics of the foundational models, such as the source of training data, the original objective of the model, training method, and evaluation metrics, among other characteristics, must be analyzed when choosing the model.
@@ 5.1.2 Preprocessing
  • Image Data: Vector embeddings were extracted from images using the Dino V2-Large model [28], a foundational computer vision model, and stored in a CSV file alongside image IDs.
@@ 5.2 Results
We compared the results with the current state of the art methods in BRSET [74, 75], and with the current state of the art foundational model RetFound [32] to predict diabetic retinopathy. The models demonstrated varying degrees of accuracy and F1 scores across different classification tasks (5-class, 3-class, and binary). The results can be seen in Table 2. Notably, the fusion model outperformed individual modality models, and current state of the art results underscoring the effectiveness of multimodal data fusion in medical diagnoses and the efficacy of our approach.
@@ 5.3.3 Preprocessing
  • Satellite Images: The embeddings of the satellite images were extracted using Variational Autoencoder (VAE) with a ResNet 50 V2 backbone (Fig. 1). This model was trained on a wider set of 81 Colombian cities’ images from 2016 to 2018. The cohort of 81 Colombian cities was chosen to expand the dataset’s diversity of landscapes by incorporating 81 cities spatially distributed throughout the country. This spatial and temporal shift allows us to avoid data leakage, incorporating at the same time a greater diversity of spatial environments that allow the model to extract better quality embeddings. Then, the embeddings were extracted from the latent space for dimensionality reduction. The VAE was used to sample all the images into a lower dimensional space, sampling to a normal distribution. The Variational Autoencoder used for processing satellite images is fundamentally based on two mathematical components: reconstruction loss and Kullback-Leibler (KL) divergence. The reconstruction loss was, in this case, the mean squared error mse, as can be seen in equation 9. The KL Divergence measures how much the learned distribution $q(z|x)$ (the encoder’s output) deviates from the prior distribution $p(z)$, which is often assumed to be a standard normal distribution. The KL divergence is given by equation 10. where $\mu$ and $\sigma$ are the mean and standard deviation of the learned distribution, and $J$ is the dimensionality of the latent space. Finally, the total loss is just the sum equation 11. In cases like satellite images, embeddings generated by foundational models such as DINO V2 may be less effective. In this case, we generated our embeddings using a VAE and compared them with those generated using DINO V2. The two methods were evaluated in predicting domestic violence. In all cases, the embeddings generated with our approach improved the results of DINO V2, as can be seen in Table 3.

### v2

@@ Abstract
Data Fusion, foundation Models, Embeddings, Multimodal Data
@@ 1 Introduction
In this era, vast amounts of data and information are generated in different settings, fields, formats, and modalities. According to the International Data Corporation (IDC), 7.5ZB are expected to be generated every year [1], and the ability to process this information and extract knowledge has never been more important [2, 3]. This data can come from diverse sources such as wearable devices, laboratory tests, Electronic Health Records (EHR), among many others, and modalities, such as medical images, clinical notes, and vital signs. The integration of data from different modalities is on the rise in Artificial Intelligence (AI) fields, such as Machine Learning (ML) and Deep Learning (DL), known as multimodal data fusion. This research area, with applications integrating areas from natural language processing to computer vision and beyond, has driven applications in healthcare [4, 5], integrating EHRs with medical images [6, 7, 8, 9, 10, 11, 12, 13] or signals from wearable devices [14, 15]. Multimodal fusion has also led to applications such as autonomous driving [16, 17], environmental sciences applications combining different sensors and satellite data[18, 19], as well as many other Internet of Things (IoT) applications and system improvements[20, 21, 22, 23, 24].
@@
Recent advances in the DL field, mainly driven by the proposal of the transformer architecture [25], have seen breakthroughs in the analysis of different modalities, such as text with models like BERT [26] or GPT [27] or images with models like Vision Transformers [28]. These models have been a revolution also in specific fields such as medicine with impressive performances in medical applications such as medical tests or diagnosis [29, 30, 31]. However, some of the limitations of working with these data modalities and these specialized DL models are, the high dimensionality of the data, the heterogeneity of the diverse data modalities and formats, and the amount of data and computational resources required to train robust models for these specific modalities [32, 33, 34]. Embeddings are low-dimensional numerical vectors learned during DL model training that preserve the original data’s most relevant features. Embeddings provide a common format that can solve data extraction and model development in environments where computational resources are scarce. Furthermore, their simplicity gives embeddings a possible way to unify the different data modalities and an efficient and generalizable solution for building multimodal models.
@@
Since embedding is proposed as a potential solution to data heterogeneity, and computational resource constraints, the DL models used to extract these embeddings must be extremely robust models trained on large amounts of data. In this context, we discuss open-source foundation models, such as Dino v2 [35] for general images or Llama 2 [36] for general text. Foundation models may also be available for specific tasks such as clinical notes with examples like Med-Bert [37], Bio-Bert [38], or medical images like Retfound for retinal images [39].
@@
However, having data and making a prediction is not enough and can cause harm if not done correctly. Adopting robust and reproducible practice is the most important characteristic of any data-related process [40]. Using AI techniques irresponsibly will lead to poor model performance in bias [41, 42, 43, 44], robustness, and fairness [45, 46, 47]. Acknowledging and addressing potential biases inherent in the data sources and models employed is pivotal.
@@
There are already previous frameworks for data fusion, starting with the best-known Joint Directors of Laboratories (JDL) [48] and its multiple updates that led to the Data Fusion Information Group (DFIG) [49] model with its most recent version that integrates AI techniques [50]. However, these models do not leverage foundation models and DL nor incorporate best practices for reducing bias and improving the model’s robustness, and have a focus only on applications such as industry and the military, leaving out use cases in environments such as healthcare.
@@
In this paper, we delve into a foundational approach for multimodal data fusion, centering the framework on the use of foundation models, vector embeddings, and good practices for data mining. Our work draws inspiration from established frameworks such as the JDL [48] and DFIG [49]. We emphasize the role of embeddings in providing flexibility and improving efficiency, robustness, and fairness, which is even more crucial in low-resource settings where errors and bias are relatively more impactful [51]. We also showed the model’s performance in three healthcare use cases: diabetic retinopathy diagnosis using fundus retinal images and EHR data, domestic violence prediction using open data from satellite images and internet data such as online news, and medical diagnosis and gender extraction using radiological images and clinical notes.
@@
  • Introduction of a Novel Data Fusion Model: We have developed an innovative process model for multimodal data fusion, rooted in the principles of foundation models and embeddings. This model is designed to efficiently address the challenges of big data’s high dimensionality and heterogeneity, making it particularly suitable for complex environments like healthcare.
@@ 2 Background
Some improvements were proposed during the following years, such as introducing adaptive generalized closed-loop data fusion processes and the perceptual reasoning system [54]. Eventually, a second revision of the JDL model (JDL II) to align with Department of Defense (DoD) priorities, improve quality control, reliability, and consistency, and empathize with the need for co-processing of inferencing processes was proposed in 2004 [55]. Also in that year, due to the shortcomings of the model, mainly related to the lack of clarity of human participation and feedback, the DFIG model was created, adding two levels (Level 5 - User Refinement, Level 6 - Mission Management) to the JDL model [40]. However, other adaptations and updates were proposed to the JDL model during the following years, mainly focusing on the optimization and reliability of the model by proposing communication across the JDL levels [56] or implementing different techniques such as game theory [57, 58], fuzzy logic [59], ontologies [60] and other mathematical modeling approaches [61, 62, 63]. Other revisions focused on the fusion, impact, and communication between hard data (machine or environmental data) and soft data (human data). Changes in specific levels such as Levels 1 and 2 incorporating the human in the fusion and preprocessing levels [64], subsequently in Levels 4 and 5 through hybrid-sensing and hybrid-cognition [65], or even the incorporating a management network in parallel to the JDL model [66, 67] were proposed.
@@
The proposed Data Fusion for Data Mining (DF-DM) model is a foundational process model for multimodal data fusion based on DFGI and CRISP-DM. The following sections will introduce the most up-to-date data fusion model (DFGI) with the adaptation for ML and AI [50], integrate the Cross-Industry Standard Process for Data Mining (CRISP-DM) process model [69], and present some modifications to implement foundation models and embeddings in the DFGI model. Finally we’ll provide three use cases to show the flexibility and adaptability of our model, as well as usability.
@@ 2.2 The Integration of the Cross-Industry Standard Process for Data Mining (CRISP-DM)
  • Many modalities, many costs: The use of multimodal data, although one of the most efficient ways to integrate and take advantage of data, also becomes very expensive when discussing high-dimensional data such as text or images. Given this, we propose using newer techniques and models for dimensionality reduction, such as extracting embeddings from foundation models or using foundation models directly for zero-shot learning tasks.
@@
In the era of AI-driven analytics, the integration of data mining and data fusion stands as a paramount challenge, primarily due to the heterogeneity of data and the constraints of computational resources. To address these challenges, we introduce an innovative methodology within our Data Fusion for Data Mining (DF-DM) model (Figure 2), which extends the existing DFGI model by integrating the model with the best practices of the CRISP-DM process model. Additionally our model leverages the use of foundation models and vector embeddings simplifies the processing of diverse data types but also significantly reduces the computational overhead, making advanced data analytics more accessible and efficient.
@@
The rationale behind the DF-DM model’s design is twofold. Firstly, by incorporating the CRISP-DM process, we aim to adhere to a proven, structured approach for data mining that emphasizes business understanding, data understanding, bias, and a cyclical process for model refinement. This ensures that our model is not only technically sound but also closely aligned with practical, real-world applications. Secondly, the decision to utilize embeddings and foundation models stems from their ability to reduce high-dimensional data into more manageable, lower-dimensional vectors. This significantly alleviates computational demands and facilitates more effective integration of diverse data types, crucial for addressing the complex challenges of multimodal data fusion in healthcare and beyond. Before delving into the methodology, it is pertinent to clarify two pivotal concepts:
  • foundation models: foundation models are at the heart of our methodology. These are extensive, pre-trained models that have been developed to capture a wide array of features across specific data modalities, such as text or images. By understanding the general characteristics inherent to each modality, these models can generate useful representations of new, unseen data. This ability is crucial for extracting relevant features without necessitating additional, computationally expensive training phases for each new dataset.
  • Embeddings: Embeddings are also a pivotal role in our methodology, transforming high-dimensional data into a more manageable, lower-dimensional space. Formally, given an input data point $x\in\mathbb{R}^{n}$, where $n$ is the dimensionality of the raw data in a high-dimensional space. We seek to find its embedding $E\in\mathbb{R}^{d}$ in a d-dimensional vector space where $d$ is the dimension of the embedding vector and $d<<n$. The embedding is generated given the transformation $E=f(x)$, where $f$ denotes the transformation function derived from the foundation model. This process significantly reduces dimensionality, with $d<<dim(X)$, facilitating more efficient data analysis.
@@ 3.1 Level 0 - Data Assessment:
  • Data preprocessing: Data preprocessing involves cleaning, transformation, normalization, encoding, reduction, and formatting, making it a critical step in the overall DF-DM model. In this step, as shown in Figure 3, using embeddings and foundation models is crucial for generating latent representations and extracting meaningful information from high-dimensional data such as images or text. Techniques such as zero-shot learning or embedding extraction are the key elements in this approach to extract features in low-resource settings. This approach significantly reduces computational costs and storage requirements and creates a unified format that facilitates multimodal data fusion.
@@ 3.2 Level 1 - Object Assessment
In Level 1, since the data received from Level 0 are embeddings or features extracted from the foundation models, the data is already in a lower dimensionality and a common format, facilitating different ML and AI techniques. As seen in Figure 3, Option 2, using embeddings, reduces the complexity and dimensionality of the data, making it easy and inexpensive to do the ML modeling because we can use simple models such as linear models. The data resulting from embedding extraction is methodically divided based on data modality (e.g., text, images). This separation enables targeted model training for each data type, enhancing specificity and accuracy. Integration of these segmented models is achieved through a comprehensive analysis framework, allowing for nuanced insights across modalities. For example, using the data from Level 0 in a CSV format due to the embedding extraction, the CSV file can be split for each modality to train a predictive model per modality. ML models can be trained to separately measure the impact of clinical notes, medical images, and tabular data in a given condition.
@@ 3.8 Level 7 - Bias Assessment
Figure 3: Illustrative Framework for Utilizing foundation models in Various Tasks. The figure assumes an initial foundation model for a general task and 3 different options. Option 1 is Zero-shot learning using the foundation model directly for a downstream task. Option 2 suggests the use of embedding, where embeddings of the original data are extracted and used for downstream task training. Option 3 means fine-tuning the full model for a specific task. The resulting model can also be used for embedding extraction.
@@ 4.2 Model Bias
Model bias can occur when the model’s architecture or training data introduces inherent bias. For example, if the foundation models used for embedding extraction have been trained on biased datasets, this bias can propagate into the DF-DM model. To address model bias, we suggest considering the following:
  • Bias Analysis of foundation models: Before embedding extraction, thoroughly analyze the foundation models to identify and mitigate potential bias. Analyze the current model available and the best option for your use case. Characteristics of the foundation models, such as the source of training data, the original objective of the model, training method, and evaluation metrics, among other characteristics, must be analyzed when choosing the model.
@@ 5.1.2 Preprocessing
  • Image Data: Vector embeddings were extracted from images using the Dino V2-Large model [35], a foundation computer vision model, and stored in a CSV file alongside image IDs.
@@ 5.2 Results
We compared the results with the current state of the art methods in BRSET [84, 85], and with the current state of the art foundation model RetFound [39] to predict diabetic retinopathy. The models demonstrated varying degrees of accuracy and F1 scores across different classification tasks (5-class, 3-class, and binary). The results can be seen in Table 2. Notably, the fusion model outperformed individual modality models, and current state of the art results underscoring the effectiveness of multimodal data fusion in medical diagnoses and the efficacy of our approach.
@@ 5.3.3 Preprocessing
  • Satellite Images: The embeddings of the satellite images were extracted using Variational Autoencoder (VAE) with a ResNet 50 V2 backbone (Fig. 1). This model was trained on a wider set of 81 Colombian cities’ images from 2016 to 2018. The cohort of 81 Colombian cities was chosen to expand the dataset’s diversity of landscapes by incorporating 81 cities spatially distributed throughout the country. This spatial and temporal shift allows us to avoid data leakage, incorporating at the same time a greater diversity of spatial environments that allow the model to extract better quality embeddings. Then, the embeddings were extracted from the latent space for dimensionality reduction. The VAE was used to sample all the images into a lower dimensional space, sampling to a normal distribution. The Variational Autoencoder used for processing satellite images is fundamentally based on two mathematical components: reconstruction loss and Kullback-Leibler (KL) divergence. The reconstruction loss was, in this case, the mean squared error mse, as can be seen in equation 9. The KL Divergence measures how much the learned distribution $q(z|x)$ (the encoder’s output) deviates from the prior distribution $p(z)$, which is often assumed to be a standard normal distribution. The KL divergence is given by equation 10. where $\mu$ and $\sigma$ are the mean and standard deviation of the learned distribution, and $J$ is the dimensionality of the latent space. Finally, the total loss is just the sum equation 11. In cases like satellite images, embeddings generated by foundation models such as DINO V2 may be less effective. In this case, we generated our embeddings using a VAE and compared them with those generated using DINO V2. The two methods were evaluated in predicting domestic violence. In all cases, the embeddings generated with our approach improved the results of DINO V2, as can be seen in Table 3.